# GFlowVLM: Enhancing Multi-step Reasoning in Vision-Language Models with Generative Flow Networks

### v1

@@ A.2 NumberLine
This task involves moving a number along a synthetic number line to reach a target. NumberLine requires identifying two numbers in an image: “Target: $x$” and “Current: $y_{t}$”, where $x$ and $y_{t}$ are both integers such that $x,y_{t}\in[n_{min},n_{max}]$. The agent’s goal is to align $y_{t}$ with $x$ by outputting an action $a_{t}$ from the discrete set {“+”, “-”, [DONE](if applicable)}. Actions “+” and “-” adjust $y_{t}\pm 1$, while [DONE] signals task completion (see detailed prompt template in Tab. 7). An episode ends when the $y_{t}=x$, or when the maximum step $T=2n_{max}$ is reached, which is the default setup of the environment. We set $n_{min}$ and $n_{max}$ as 0 and 5, respectively for the in-distribution examples, and set $n_{min}$ and $n_{max}$ as 10 and 50 for generating OOD examples. In the reward function used in [41], an agent receives a reward of $r(s_{t},a_{t})=1$ when $y_{t+1}=x$, a penalty of $r(s_{t},a_{t})=-1$ upon taking an action that does not move the current number $y_{t}$ to the target $x$, and a reward of $r(s_{t},a_{t})=0$, otherwise. For GFlowVLM, we revise the reward function with non-negative values as follows:
$R(x,y_{t})=\frac{c}{|x-y_{t}|+1}$ (9)
@@
where $c$ is a scaling constant set to 100. This reward incentivizes the model to bring the current number closer to the target, progressively increasing the reward as the gap decreases. For fair comparison, we run RL4VLM [41] with revised reward structure.
A.3 Blackjack
@@
The Blackjack task requires the VLM to reason with visual information and adapt to stochastic outcomes. The observation $o_{t}$ includes two dealer cards (one face-down) and the player’s cards. The agent aims to win by selecting an action $a_{t}$ from {“stand”, “hit”, [DONE](if applicable)} (see detailed prompt template in  Tab. 8). In the reward function used in [41], an agent receives a reward of $r(s_{t},a_{t})=1,0,-1$ upon win, draw and loss, respectively. We revise the reward function to replace non-negative values as follows:
$R(s_{t},a_{t})=\max(1\times 10^{-10},(r(s_{t},a_{t})+1)\times 10),$ (10)
@@
where $r(s_{t},a_{t})$ represents the environment’s original reward for state $s_{t}$ and action $a_{t}$. This scales the rewards and ensures they are strictly non-negative. For fair comparison, we run RL4VLM [41] with revised reward structure.
Appendix B Training Objectives
@@
B.1 Variance Trajectory Balanced (Var-TB) Loss
The Trajectory-Balanced (TB) objective [23] ensures that the probability of generating a complete trajectory $\tau=(s_{0}\rightarrow s_{1}\rightarrow\cdots\rightarrow s_{n}\rightarrow x)$ is proportional to the reward $R(\tau)$. Under the Markovian assumption, the forward policy $P_{F}(s_{t}|s_{t-1})$ transitions from state $s_{t-1}$ to $s_{t}$, while the backward policy $P_{B}(s_{t-1}|s_{t})$ ensures consistency between forward and backward flows. This objective is given by:
@@ B.1 Variance Trajectory Balanced (Var-TB) Loss
Under the non-Markovian assumption of generating a complete trajectory $\tau=(z_{0}\rightarrow z_{1}\rightarrow\cdots\rightarrow z_{n}\rightarrow x)$, and after adding goal into condition, we have:
@@
$\begin{split}\mathcal{L}_{\mathrm{VarTB}}(\boldsymbol{\tau};\theta)=\frac{1}{N}\sum_{i=1}^{N}\bigg{(}\mathrm{\zeta}(\tau_{i};\theta,g)-\mathbb{E}_{\boldsymbol{\tau}}\big{[}\mathrm{\zeta}(\boldsymbol{\tau};\theta,g)\big{]}\bigg{)}^{2},\end{split}$ (14)
where $N$ represents the number of sampled trajectories. This loss ensures that high-reward trajectories are sampled more frequently by the policy.
B.2 Subtrajectory Balanced (SubTB) Loss
@@
$\begin{split}\mathcal{L}_{\mathrm{SubTB}}(z_{0:m},g;\theta)=\sum_{0\leq i<j\leq m}\\ \Bigg{(}\log\frac{R(z_{0:i}\top)\prod_{k=i+1}^{j}P_{F}(z_{k}|z_{0:k-1},g;\theta)P_{F}(\top|z_{0:j},g;\theta)}{R(z_{0:i}\top)P_{F}(\top|z_{0:i},g;\theta)}\bigg{)}^{2}\end{split}$ (16)
where $\top$ is the [DONE] symbol, denoting the terminal state, and process continues until [DONE] symbol $\top$ is generated similar to [8]. This loss penalizes discrepancies in local transitions and ensures that all subsegments of a trajectory follow the correct balance conditions, reducing variance in smaller parts of the trajectory.
B.2.1 Detailed Balanced (DB) Loss
@@
Appendix C Details of Experimental Setup
@@
C.1 Off-Policy Data Collection
@@ C.3 Diversity Metric
C.4 General Setup for Baselines and GFlowVLM
@@
C.5 CoT Weighting Factor $\lambda$
@@
Appendix D Qualitative Results
@@
Appendix E Ablation Study of Markovian and Non-Markovian

### v2

@@ B.2 NumberLine
This task involves moving a number along a synthetic number line to reach a target. NumberLine requires identifying two numbers in an image: “target: $c$” and “current: $y_{t}$”, where $c$ and $y_{t}$ are both integers such that $c,y_{t}\in[n_{min},n_{max}]$. The agent’s goal is to align $y_{t}$ with $c$ by outputting an action $a_{t}$ from the discrete set {“+”, “-”, [DONE](if applicable)}. Actions “+” and “-” adjust $y_{t}\pm 1$, while [DONE] signals task completion (see detailed prompt template in Tab. 7). An episode ends when the $y_{t}=x$, or when the maximum step $T=2n_{max}$ is reached, which is the default setup of the environment. We set $n_{min}$ and $n_{max}$ as 0 and 5, respectively for the in-distribution examples, and set $n_{min}$ and $n_{max}$ as 10 and 50 for generating OOD examples. In the reward function used in [41], an agent receives a reward of $r(s_{t},a_{t})=1$ when $y_{t+1}=c$, a penalty of $r(s_{t},a_{t})=-1$ upon taking an action that does not move the current number $y_{t}$ to the target $c$, and a reward of $r(s_{t},a_{t})=0$, otherwise. For GFlowVLM, we revise the reward function with non-negative values as GFlowNets inherently require non-negative as follows:
$R(x)=R(c,y_{t})=\frac{l}{|c-y_{t}|+1}$ (9)
@@
where $l$ is a scaling constant set to 100. This reward incentivizes the model to bring the current number closer to the target, progressively increasing the reward as the gap decreases. For fair comparison, we run RL4VLM [41] with revised reward structure.
B.3 Blackjack
@@
The Blackjack task requires the VLM to reason with visual information and adapt to stochastic outcomes. The observation $o_{t}$ includes two dealer cards (one face-down) and the player’s cards. The agent aims to win by selecting an action $a_{t}$ from {“stand”, “hit”, [DONE](if applicable)} (see detailed prompt template in  Tab. 8). In the reward function used in [41], an agent receives a reward of $r(x)=1,0,-1$ upon win, draw and loss, respectively. Since GFlowNets inherently require non-negative reward, we revise the reward function to replace non-negative values as follows:
$R(x)=\max(1\times 10^{-10},(r(x)+1)\times 10),$ (10)
@@
where $r(x)$ represents the environment’s original reward for state $x$. This scales the rewards and ensures they are strictly non-negative. For fair comparison, we run RL4VLM [41] with revised reward structure.
$R(x)$ represents the desirability or quality of a complete trajectory with final state $x$, similar to RL. It defines the target distribution from which the GFlowNet learns to sample, where higher-reward outcomes should be sampled more frequently.
Appendix C Training Objectives
@@
C.1 Variance Trajectory Balanced (Var-TB) Loss
The Trajectory-Balanced (TB) objective [23] ensures that the probability of generating a complete trajectory $\tau=(s_{0}\rightarrow s_{1}\rightarrow\cdots\rightarrow s_{n}=x)$ is proportional to the reward $R(\tau)$. Under the Markovian assumption, the forward policy $P_{F}(s_{t}|s_{t-1})$ transitions from state $s_{t-1}$ to $s_{t}$, while the backward policy $P_{B}(s_{t-1}|s_{t})$ ensures consistency between forward and backward flows. This objective is given by:
@@ C.1 Variance Trajectory Balanced (Var-TB) Loss
Under the non-Markovian assumption of generating a complete trajectory $\tau=(z_{0}\rightarrow z_{1}\rightarrow\cdots\rightarrow z_{n}=x)$, and after adding goal into condition, we have:
@@
$\begin{split}\mathcal{L}_{\mathrm{VarTB}}(\boldsymbol{\tau};\theta)=\frac{1}{K}\sum_{k=1}^{K}\bigg{(}\mathrm{\zeta}(\tau_{k};\theta,g)-\mathbb{E}_{\boldsymbol{\tau}}\big{[}\mathrm{\zeta}(\boldsymbol{\tau};\theta,g)\big{]}\bigg{)}^{2},\end{split}$ (14)
where $K$ represents the number of sampled trajectories. This loss ensures that high-reward trajectories are sampled more frequently by the policy.
C.2 Subtrajectory Balanced (SubTB) Loss
@@
$\begin{split}\mathcal{L}_{\mathrm{SubTB}}(z_{0:m},g;\theta)=\sum_{0\leq i<j\leq m}\\ \Bigg{(}\log\frac{R(z_{0:i}\top)\prod_{k=i+1}^{j}P_{F}(z_{k}|z_{0:k-1},g;\theta)P_{F}(\top|z_{0:j},g;\theta)}{R(z_{0:j}\top)P_{F}(\top|z_{0:i},g;\theta)}\bigg{)}^{2}\end{split}$ (16)
where $\top$ is the [DONE] symbol, denoting a trivial terminal state, and process continues until [DONE] symbol $\top$ is generated similar to [8]. This loss penalizes discrepancies in local transitions and ensures that all subsegments of a trajectory follow the correct balance conditions, reducing variance in smaller parts of the trajectory.
C.3 Detailed Balanced (DB) Loss
@@
Comparisons of Loss Functions
TB loss controls the variance of $\zeta$ for the sampled trajectories, not the individual trajectory. Its main role is to bias sampling so that trajectory selection probability aligns with rewards [9]. In addition, DB loss excels with dense rewards by ensuring flow consistency at each state, while SubTB and TB perform better in sparse settings by optimizing flow across (sub)trajectories. Additionally, TB is suited for tasks with known full sequences, and SubTB for costly large-trajectory sampling.
Computational Complexity
In practice, we calculate (sub)trajectory or transition-based loss functions, which operate over (sub)trajectories or sampled transitions rather than the full state space. This allows us to efficiently handle the non-Markovian dependencies with linear complexity.
Appendix D Details of Experimental Setup
@@
D.1 Off-Policy Data Collection
@@ D.3 Diversity Metric
D.4 General Setup for Baselines and GFlowVLM
@@
D.5 CoT Weighting Factor $\lambda$
@@
Appendix E Qualitative Results
@@
Appendix F Ablation Study of Markovian and non-Markovian